OpenStudy (anonymous):

The Navstar Global Positioning System (GPS) utilizes a group of 24 satellites orbiting the Earth. Using "triangulation" and signals transmitted by these satellites, the position of a receiver on the Earth can be determined to within an accuracy of a few centimeters. The satellite orbits are distributed around the Earth, allowing continuous navigational "fixes." The satellites orbit at an altitude of approximately 11,000 nautical miles [1 nautical mile = 1.852 km=6076ft]. Part A Determine the speed of each satellite. Part B Determine the period of each satellite.

OpenStudy (anonymous):

@agent0smith I need help with this last one too,I am doing something wrong again this is what I

OpenStudy (anonymous):

did

OpenStudy (anonymous):

m v^2 / r = G M m / r^2 v = sqroot of GM/r v= sqroot of (6.67 x 10^-11) (5.98 x 10^24) / (6381852) v = 7910 m/s

OpenStudy (anonymous):

but for some reason this is wrong

OpenStudy (anonymous):

for the radius i added 1852 + 6380000 = 6381852 m

OpenStudy (loser66):

1852km = 1.852 x 10^6m + something, . why do you just let it there without convert to m?

OpenStudy (anonymous):

I have to convert bc G constant is in meters

OpenStudy (loser66):

sure!! but the altitude of the satellite is 1852 km, right?

OpenStudy (anonymous):

and the velocity has to be in meters

OpenStudy (anonymous):

it says its 1.852 km so that is 1852 m

OpenStudy (loser66):

so, the radius = 1852 x 10^6 + radius of earth ,

OpenStudy (anonymous):

i dont understand where you got the 10^6 from..why would i multiply 1852 by 10^6 ?

OpenStudy (loser66):

ok, I misread

OpenStudy (loser66):

but the altitude is 11000 nautical and 1 nautical = 1.852km so the altitude is 11000*1852 to get m from ALTITUDE only, right ? then r = that altitude + radius of Earth

OpenStudy (anonymous):

ahhh, i see it, i dont know but I will try it out and see if it works

OpenStudy (anonymous):

sorry it took long for some reason i lost connection to openstudy finally its back open...I got an answer of 3.86 x 10^14 m/s ...does that seem right?

OpenStudy (anonymous):

i get an answer of 3.86 x 10^14 m/s does that seem right?

OpenStudy (anonymous):

never mind i just checked again on the calculator i think its 3861. 32 m/s

OpenStudy (anonymous):

thank you @Loser66

OpenStudy (anonymous):

thank you so much for your help @agent0smith and @Loser66

OpenStudy (agent0smith):

3.86 x 10^14 m/s this is about a million times faster than the speed of light, which should've been a clue that your answer was a bit off :P

OpenStudy (anonymous):

i realized after a while that seemed wrong, thank you